# Best Agent Identification for General Game Playing

### v1

@@ 4.1.1 Number of rounds
One advantage of our proposed algorithm over alternatives such as Successive Reject, is that the total number of rounds $n$ has no impact on which arms are pulled each round. This means that the predicted best arm for each bandit $J_{m}$ can be returned after any number of rounds, without a loss in prediction performance. This “stop anytime” nature also allows for alternative stopping criteria, such as algorithm runtime or prediction confidence, to be easily enforced. It is also possible to add additional rounds (i.e., increase the value of $n$) from an arbitrary stopping point, again without any loss in performance.
4.1.2 Alpha value
@@
The alpha value used for our Wilson score interval determines the desired level of confidence, where a lower alpha value typically equates with a higher degree of exploration (vice versa). In line with the typical exploration-exploitation trade-off, higher alpha values often perform better in the short term and lower alpha values perform better in the long term.
5 Experiments
5.1 Game-Agent Datasets
While the majority of prior research on best arm identification has focused on proving certain theoretical results for specific algorithms, we will instead demonstrate the practical application and empirical results of our approach on real-world data. Specifically, we will use results obtained from the GVGAI and Ludii domains across a variety of games and agents.
Results for GVGAI were sourced from the work of Stephenson et al. (2020), providing a total of 3,990,760 trials across 108 games and 27 agents. These trials are also split amongst the five distinct levels available for each game, allowing us to include agent results from a subset of levels if desired. To evaluate whether agent performance differences between levels affects our approach’s effectiveness, we define both a GVGAI (all levels) dataset that includes trials for each game across all five levels, as well as a GVGAI (level 1) dataset that includes only the 798,516 trials from the first level of each game.
Results for Ludii were obtained by running 29 variants of an Alpha Beta agent, each using a different heuristic state evaluation function (Stephenson et al., 2021). The Alpha Beta agent was set to Player 1, while any other players were controlled by a standard Upper Confidence bounds applied to Trees (UCT) algorithm (Kocsis and Szepesvári, 2006). Each agent was given 0.5 seconds of thinking time per move. A maximum turn limit of 500 was applied to all games, and exceeding this value results in a draw for all remaining active players. Each agent (i.e., Alpha Beta heuristic) was run for 100 trials on each of the 1085 games included in Ludii v1.3.9, giving a total of 3,434,627 trials. Results were generated using a single general node (AMD EPYC 7551 @2.55Ghz, 256GB DDR4 @2666Mhz) of the DeepThought high-performance computer (Flinders University, 2021).
For our presented experiments, rather than running new trials for a chosen game-agent pair we will instead sample rewards randomly from these datasets of previously performed trials. The number of pre-calculated results included in these datasets is large enough to act as a proxy for the true reward distribution of each game-agent pair, and allows for both an increased number of repeated comparison runs and easily reproducible results. Fixing our result datasets also provides us with a defined “ground truth” measure of agent performance, which is necessary for the simple regret calculations that will be used to evaluate and compare different best-agent identification algorithms.
All source code, game-agent datasets and regret scores used to produce the presented results is publicly available online. [^1] https://github.com/stepmat/Best_Agent_Identification_GGP
@@ 5.1 Game-Agent Datasets
GapE Selects arms that are close to the current best arm in each bandit Exploration value
$a$ = [1, 2, 4, 8, 16]
GapE-V Alternative version of GapE that also considers reward variance Exploration value
$a$ = [1, 2, 4, 8, 16]
UCB-E A highly exploring version of UCB that selects arms with the highest upper confidence bound ($a=2\cdot\log(n)$ is equivalent to UCB1) Exploration value
$a$ = [2, 4, 8, 16] $\cdot$ $\log(n)$, where $n$ = round number
@@
Optimistic-WS Selects arms with the highest regret change potential, based on Wilson score estimates Exploration value
$c$ = [1, 2, 4, 8, 16]
@@
Table 1: Best Arm Identification algorithms to compare.
5.2 Algorithm Comparison
@@
Using the GVGAI and Ludii game-agent results datasets described above, we evaluate the performance of the multi-bandit best arm identification algorithms described in Table 1. This includes our proposed Optimistic-WS algorithm, along with all previous alternatives described in Section 3.3. To provide a more standard performance baseline, we also evaluate a Random sampling approach that selects a bandit-arm pair at random each time with equal probability (i.e., with replacement), and a Uniform sampling approach that selects each bandit-arm pair an equal number of times (i.e., without replacement). For algorithms with adjustable hyperparameters, a range of suitable values were tested. Each algorithm and/or hyperparameter value was run on both datasets 10 times, with 50,000 rounds (i.e., arm pulls) in each run. The simple regret of each algorithm was calculated every 1000 pulls, providing us with an average measure of performance across a varying number of rounds.
@@
To prevent cluttering in our presented results, we have selected only the overall best performing hyperparameter value for each algorithm to compare. More specifically, for GapE ($a$ = 2), for GapE-V ($a$ = 1), for UCB-E ($a$ = 2), and for Optimistic-WS ($a$ = 16). The decision of which hyperparameter value was the “best performing” was determined by whichever value had the lowest average simple regret for the highest number of rounds. In practice, this choice of hyperparameter value had little impact on the effectiveness of our Optimistic-WS algorithm, which always managed to outperform all other approaches. Full results for all hyperparameter values are provided in the Appendix.
5.3 Results
5.3.1 GVGAI
@@
The average regret of each best agent identification algorithm for both the GVGAI (all levels) and GVGAI (level 1) datasets are shown in Figures 1 and 2 respectively. From these results we can see that our proposed Optimistic-WS algorithm provides a substantial performance improvement compared to all previous approaches for any number of rounds. When applied to the GVGAI (all levels) dataset, the simple regret of our Optimistic-WS approach is approximately 35.5% smaller than that of the second best approach (UCB-E), and 67.9% smaller than the more typical uniform sampling approach (averaged across the full 50,000 rounds). The improvement provided by Optimistic-WS is even more pronounced for the GVGAI (level 1) dataset, with our simple regret being 47.0% smaller than UCB-E and 70.4% smaller than uniform sampling over the same 50,000 rounds period.
@@
5.3.2 Ludii
The average regret of each best agent identification algorithm, for the Ludii game-agent dataset is shown in Figure 3. Similar to the GVGAI datasets, our proposed Optimistic-WS algorithm again provides a significant performance increase, achieving a simple regret that is 35.1% smaller than UCB-E and 55.8% smaller than uniform sampling when averaged across all 50,000 rounds.
One immediate observation is that the reduction in average simple regret for the Ludii dataset is less rapid compared to the GVGAI dataset. This could be due to several factors, but is most likely caused by the fact that the Ludii domain simply has more games compared to GVGAI, and hence more bandits to split pulls between. As a result, we would require a much larger number of rounds before our regret would begin to plateau, although we can see this beginning to occur for Optimistic-WS towards the 50,000 round mark.
@@ 5.3.2 Ludii
5.4 Discussion
Looking first at the GVGAI dataset results, see Figures 1 and 2, we can identify some general trends for each algorithm. Interestingly, the performance comparisons between the GVGAI (all levels) and GVGAI (level 1) datasets are remarkably similar, and so our observations can be jointly applied to both.
Optimistic-WS clearly performs best, achieving a lower average simple regret than all other algorithms for any number of rounds. After this, the second best performing approach is most often UCB-E, with the default UCB1 exploration value of $a=2\cdot\log(n)$. However, after this the comparison is not as clear cut, with different algorithms performing better after varying round amounts. Gape-E starts off well early on but performs worse as the number of rounds increases, eventually only beating Random and Uniform after 50,000 rounds. GapE-V by comparison starts off worse, but overtakes GapE around the 15,000 round mark. Both Successive Rejects and Sequential Halving are designed for a fixed total number of arm pulls, and so initially perform very poorly as the are not intended for early stopping. However, they eventually match the performance of other top performing models, such as UCB-E, once we reach defined 50,000 round end point. The Anytime Sequential Halving algorithm demonstrates its improved “stop anytime” functionality over regular Sequential Halving, achieving a much lower average simple regret in the earlier rounds and a near equal performance after 50,000 rounds. Lastly, it is unsurprising that Random and Uniform performed the worst overall, both of which are fairly naive (although still commonly used) evaluation approaches.
Looking next at the Ludii dataset results, see Figure 3, we can see that Optimistic-WS once again significantly outperforms all other approaches. UCB-E started off well, but eventually worsened and ended up matching the performance of GapE after 50,000 rounds). GapE-V also appeared to perform worse than its GapE counterpart, as did Anytime Sequential Halving. This apparent improvement in GapE compared to the GVGAI results could be due to the fixed 50,000 round limit, and that with additional rounds the relative performance of GapE may worsen (although this is unconfirmed). Lastly, the results for Successive Rejects and Sequential Halving are particularly odd, flattening well before the 50,000 round limit. The reason for this is largely due to their requirement of a fixed number of arm pulls per bandit, and is discussed in greater detail below.
@@ 5.4 Discussion
One significant drawback of the Sequential Halving and Successive Reject algorithms is that they often do not use all of the pulls available to them (Audibert et al., 2010; Karnin et al., 2013). This is because both algorithms distribute pulls across several arm selection rounds ahead of time, requiring that the number of pulls each round be divisible between the remaining arms, which often results in some leftover pulls. This problem is particularly prevalent in situations where the number of available pulls is relatively low compared to the number of bandit arms. Our Ludii dataset presents just such a scenario, with 1085 games and a budget of 50,000 trials meaning that each game is assigned only 46 trials. Given that each game can have up to 29 potential agents that need evaluating, this very small trials-to-agents ratio severely limits the effectiveness of these approaches. Sequential Halving is most impacted by this, with only 26,834 ($\sim$54%) of its allocated trials being used, resulting in a flat regret line after this point. Successive Rejects is also affected by this same issue, using just 39,102 ($\sim$78%) of its available trials, but the fact that it discards its arms one at a time instead of removing half appears to reduce this impact. This issue is also much less prevalent for the GVGAI dataset, likely due to the reduced number of games this dataset has, leading to an increased number of trials per game-agent pair.

### v2

@@ IV-A1 Number of rounds
One advantage of our proposed algorithm over alternatives such as Successive Rejects, is that the total number of rounds $n$ has no impact on which arms are pulled each round. This means that the predicted best arm for each bandit $J_{m}$ can be returned after any number of rounds, without a loss in prediction performance. This “stop anytime” nature also allows for alternative stopping criteria, such as algorithm runtime or prediction confidence, to be easily enforced. It is also possible to add additional rounds (i.e., increase the value of $n$) from an arbitrary stopping point, again without any loss in performance.
IV-A2 alpha value
@@
The alpha value directly modifies the upper and lower bounds returned by our chosen confidence interval. A lower alpha value results in a higher degree of ”optimism/pessimism” in the potential regret change of each arm, typically leading to an increase in overall arm exploration (vice versa). In line with the typical exploration-exploitation trade-off, higher alpha values often perform better in the short term and lower alpha values tend to perform better in the long term.
IV-B Uniform Bandit Selection Variant
One potential variant of the RCP algorithm is to select bandits uniformly and then apply the RCP algorithm within each bandit individually. This variation essentially treats our multi-bandit domain as a set of independent best-arm identification problems, where each bandit gets an equal number of pulls. This comes with a potential downside in that we can no longer redistribute arm pulls between bandits (i.e., if we are already confident in the best arm for one bandit, we cannot reallocate it’s remaining arm pulls to other bandits). However, it is also possible that this approach could lead to better performance in cases where our RCP approach either over-samples a bandit due to high outcome variance (e.g., stochastic games) or under-samples it due to unlucky outcomes causing a top-performing arm to be dismissed prematurely.
TABLE I: Best arm identification algorithms and hyperparameter values to be compared.
@@ IV-B Uniform Bandit Selection Variant
GapE Selects arms that are close to the current best arm in each bandit Exploration value ($a$) = [1, 2, 4, 8, 16]
GapE-V Alternative version of GapE that also considers reward variance Exploration value ($a$) = [1, 2, 4, 8, 16]
UCB-E A highly exploring version of UCB that selects arms with the highest upper confidence bound ($a=2\cdot\log(n)$ is equivalent to UCB1) Exploration value ($a$) = [2, 4, 8, 16] $\cdot$ $\log(n)$, where $n$ = round number
@@
RCP Selects arms with the highest regret change potential, based on a selected confidence interval estimate Alpha value ($\alpha$) = [0.1, 0.05, 0.01]
Confidence interval = [Wilson Score, Jeffreys, Clopper–Pearson, Agresti–Coull, KL-UCB, PAC]
@@
V Experiments
V-A Game-Agent Datasets
While the majority of prior research on best arm identification has focused on proving certain theoretical results for specific algorithms, we instead demonstrate the practical application and empirical results of our approach on real-world data. Specifically, we use results obtained from both the GVGAI and Ludii domains across a variety of games and agents.
Results for GVGAI were sourced from the work of [undefar], providing a total of 3,988,557 trials across 108 games and 27 agents. These trials are also split amongst the five distinct levels available for each game, allowing us to include agent results from a subset of levels if desired. To evaluate whether agent performance differences between levels affects our approach’s effectiveness, we define both a GVGAI (all levels) dataset that includes trials for each game across all five levels, as well as a GVGAI (level 1) dataset that includes only the 798,168 trials from the first level of each game.
Results for Ludii were obtained by running 29 variants of an Alpha-Beta agent, each using a different heuristic state evaluation function [undefat]. The Alpha-Beta agent was set to Player 1, while any other players were controlled by a standard Upper Confidence bounds applied to Trees (UCT) algorithm [undefac]. Each agent was given 0.5 seconds of thinking time per move. A maximum turn limit of 500 was applied to all games, and exceeding this value results in a draw for all remaining active players. Each agent (i.e., Alpha-Beta heuristic) was run for 100 trials on each of the 1085 games included in Ludii v1.3.9, giving a total of 3,434,627 trials. Any agent heuristics that were incompatible with certain game types were excluded, leading to a varying number of valid agent trials for each game. Results were generated using a single general node (AMD EPYC 7551 @2.55Ghz, 256GB DDR4 @2666Mhz) of the DeepThought high-performance computer [undefo].
For our presented experiments, rather than running new trials for a chosen game-agent pair we instead sampled rewards randomly (with replacement) from these datasets of previously performed trials. The number of pre-calculated results included in these datasets is large enough to act as a proxy for the true reward distribution of each game-agent pair, and allows for both an increased number of repeated comparison runs and easily reproducible results. Fixing our results datasets also provides us with a defined “ground truth” measure of agent performance, which is necessary for the simple regret calculations that are used to evaluate and compare different best-agent identification algorithms.
In terms of algorithm runtime, most approaches can be run reasonably fast by understanding that the confidence intervals for each game-agent pair are independent of both each other and of the current number of rounds. This means that only the upper and lower confidence bounds of the selected game-agent pair need to be recomputed, and all others will remain unchanged (essentially meaning that just one new calculation needs to be performed each round). The only exceptions to this are the KL-UCB and PAC confidence intervals, which utilise the total number of samples (i.e., current round number) in their calculations. For these confidence intervals, the upper/lower bounds of every game-agent pair must be updated after each round, essentially resulting in $MK$ calculations. However, even for the relatively large number of game-agent pairs in the Ludii dataset, these calculations still take significantly less time than it would take to run a full trial for most games. Given that the total runtime of each algorithm depends heavily on utilised hardware and software optimisations, it has been excluded from the presented results.
All source code, game-agent datasets and performance scores used to produce the presented results is publicly available online. [^1] https://github.com/stepmat/Best_Agent_Identification_GGP
V-B Dataset Exploration
Looking closer at both the GVGAI and Ludii results datasets reveals many of the domain specific factors previously mentioned in Sections II-B and III-D.
Figure 1: Distribution of win-rates for each game-agent pair in the GVGAI (all levels) dataset (total of 2913 game-agent pairs).
Figure 2: Distribution of win-rates for each game-agent pair in the Ludii dataset (total of 23,675 game-agent pairs).
  • First, the result distributions for each game-agent pair are often highly skewed towards the 0 and 1 extremities, see Figures 1 and 2, as most agents are reasonably consistent in either winning or losing a game over repeated trials. For the Ludii dataset, we can also observe an additional, albeit smaller, spike around the 0.5 win-rate for games that are frequently drawn.
  • Second, while the majority of games have only a single ”best agent”, there are a significant number of games where multiple agents achieve the same result. For the GVGAI dataset, 26 of the 108 games have multiple top-performing agents ($\sim$24%), whereas for the Ludii dataset this rises to 370 of the 1085 games ($\sim$34%). For the purpose of calculating our average probability of error results, selecting any one of these top performing agents would be considered as the optimal choice.
  • Lastly, all game trials in the GVGAI dataset have binary outcomes (0 or 1) whereas approximately 22% of game trials in the Ludii dataset have a non-binary outcome (i.e., a draw or partial win/loss result). This shows that while our Ludii results are not strictly binomial, the majority of game trials ($\sim$78%) do return either a 0 or 1 outcome.
V-C Algorithm Comparison
@@
Using the GVGAI and Ludii game-agent results datasets described above, we evaluate the performance of the multi-bandit best arm identification algorithms described in Table I. This includes our proposed RCP algorithm, using each of the six confidence intervals described in section III-D, along with all previous state-of-the-art alternatives described in Section III-C. To provide a more standard performance baseline, we also evaluate a Random sampling approach that selects a bandit-arm pair at random each time with equal probability (i.e., with replacement) and a Uniform sampling approach that selects each bandit-arm pair an equal number of times (i.e., without replacement). For algorithms with adjustable hyperparameters, a range of suitable values were tested. Each algorithm and hyperparameter value combination was run on both datasets 10 times, with 50,000 rounds (i.e., arm pulls) in each run. The simple regret of each algorithm was calculated every 1000 pulls, providing us with an average measure of performance across a varying number of rounds.
@@
VI Results
VI-A Hyperparameter Values
We first discuss the impact that varying the hyperparameter values for each best arm identification algorithm had on it’s performance. To avoid cluttering this section with a high number of graphs, the results of this comparison have been placed in the Appendices A, B and C.
VI-A1 RCP: Alpha Value
Figures 16, 17 18 (GVGAI all levels), 22, 23, 24 (GVGAI level 1), and 28, 29, 30 (Ludii) show the results of RCP for each confidence interval estimate, using alpha values 0.1, 0.05 and 0.01. From these, we can see that altering the alpha value does not substantially affect the overall trends and performance of these confidence intervals for our RCP algorithm. While there are some minor visible deviations between the alpha values, it is not possible to separate this from normal result variation (i.e., noise) and it would appear that the RCP algorithm performs near equivalently for reasonable alpha values across all of our general game playing domains.
VI-A2 RCP: Confidence Interval
Given that the performance of each RCP confidence interval estimate was largely unaffected by the chosen alpha value, we will focus specifically on the $\alpha$ = 0.05 case (Figures 17, 23 and 29). We also see very similar trends across each dataset, and so will discuss the relative performance of each RCP confidence interval estimate collectively across both general game playing domains.
First, we can see that the performance difference in terms of average simple regret for each of the first four confidence intervals (Wilson Score, Jeffreys, Clopper–Pearson and Agresti–Coull) is very minimal. All of these estimates demonstrate near identical trends in terms of regret reduction and variance, indicating that any of these confidence intervals would provide a suitable upper and lower regret bound.
In contrast, KL-UCB initially performed equally well in the early rounds, but eventually diverged from the other approaches. Interestingly, for the GVGAI datasets (all levels and level 1) KL-UCB showed significantly worse performance between 5,000 and 10,000 rounds, but was able to recover to near equal regret reduction soon after this. For the Ludii dataset, KL-UCB started to diverge from the other top performing estimates after the 10,000 round mark, and this continued to widen as more rounds were carried out.
Lastly, PAC achieved the worst overall performance, achieving a lower average simple regret than the other confidence interval estimates after any number of rounds. Despite this, PAC is the only approach to utilise an adaptive alpha value that ensures widening confidence intervals for each arm as the number of rounds increases (so long as it is not pulled) providing us with a fixed-confidence guarantee. In practice however, it seems that the four binomial proportion confidence intervals implemented (Wilson Score, Jeffreys, Clopper–Pearson and Agresti–Coull) achieve the best regret reduction results up to the 50,000 rounds evaluated.
VI-A3 GapE / GapE-V / UCB-E: Exploration Value
Figures 13, 14, 15 (GVGAI all levels), 19, 20, 21 (GVGAI level 1), and 25, 26, 27 (Ludii) show results for the GapE, GapE-V and UCB-E algorithms using a range of exploration values. From these, we can see that larger exploration values are typically associated with a higher average regret, particularly after the early rounds. The only exception to this seems to be GapE, where an exploration value of aa=2 appears to perform slightly better than aa=1.
VI-A4 Summary
Based on these findings, we have selected only the overall best performing hyperparameter value combination for each algorithm to directly compare against each other. More specifically, for GapE ($a$ = 2), for GapE-V ($a$ = 1), for UCB-E ($a$ = 2), and for RCP ($\alpha$ = 0.05, confidence interval = Wilson Score). In practice, we observed that the choice of hyperparameter values for our RCP algorithm had very little impact on its overall performance (with the exception of KL-UCB and PAC confidence intervals). The decision to only compare these specific hyperparameter combinations is simply to prevent graph overcrowding in our presented results. Full results for all hyperparameter values are provided in the Appendix.
VI-B GVGAI
@@
The average regret of each best agent identification algorithm for both the GVGAI (all levels) and GVGAI (level 1) datasets are shown in Figures 3 and 4 respectively. From these results we can see that our proposed RCP algorithm provides a substantial performance improvement compared to all previous approaches for any number of rounds. When applied to the GVGAI (all levels) dataset, the simple regret of our RCP approach is approximately 35.7% smaller than that of the second best approach (UCB-E), and 64.5% smaller than the more typical uniform sampling approach (averaged across the full 50,000 rounds). The improvement provided by RCP is even more pronounced for the GVGAI (level 1) dataset, with our simple regret being 47.3% smaller than UCB-E and 70.5% smaller than uniform sampling over the same 50,000 round period.
Looking closer at these results, we can identify some general trends for each algorithm. Interestingly, the performance comparisons between the GVGAI (all levels) and GVGAI (level 1) datasets are remarkably similar, and so our observations can be jointly applied to both. RCP clearly performs best, achieving a lower average simple regret than all other algorithms for any number of rounds. After this, the second best performing approach is most often UCB-E, with the default UCB1 exploration value of $a=2\cdot\log(n)$. However, after this the comparison is not as clear cut, with different algorithms performing better after varying round amounts. Gape-E starts off well early on but performs worse as the number of rounds increases, eventually only beating Random, Uniform and Sequential Halving after 50,000 rounds. GapE-V by comparison starts off worse, but overtakes GapE after the 20,000 round mark. Both Successive Rejects and Sequential Halving are designed for a fixed total number of arm pulls, and so initially perform very poorly as the are not intended for early stopping. However, they eventually match the performance of most other approaches once we reach defined 50,000 round end point. The Anytime Sequential Halving algorithm demonstrates its improved “stop anytime” functionality over regular Sequential Halving, achieving a much lower average simple regret in the earlier rounds and a near equal performance after 50,000 rounds. Lastly, it is unsurprising that Random and Uniform performed the worst overall, both of which are fairly naive (although still commonly used) evaluation approaches.
@@
VI-C Ludii
The average regret of each best agent identification algorithm for the Ludii game-agent dataset is shown in Figure 5. Similar to the GVGAI datasets, our proposed RCP algorithm again provides a significant performance increase, achieving a simple regret that is 35.5% smaller than UCB-E and 55.7% smaller than uniform sampling when averaged across all 50,000 rounds.
One immediate observation is that the reduction in average simple regret for the Ludii dataset is less rapid compared to the GVGAI dataset. This could be due to several factors, but is most likely caused by the fact that the Ludii domain simply has more games compared to GVGAI, and hence more bandits to split pulls between. As a result, we would require a much larger number of rounds before our regret would begin to plateau, although we can see this beginning to occur for RCP towards the 50,000 round mark.
In terms of the other other algorithms evaluated, UCB-E started off well (even outperforming RCP for the first few thousand rounds) but quickly worsened and ended up matching the performance of GapE after the full 50,000 rounds. This apparent improvement in GapE compared to the GVGAI results could be due to 50,000 round limit being relatively small compared to the number of Ludii game-agent pairs (23,675) and that with additional rounds the relative performance of GapE may worsen (although this is unconfirmed). GapE-V appeared to perform consistently worse than its GapE counterpart, perhaps due to an insufficient number of rounds for it to get an accurate estimation of each arm’s variance. Anytime Sequential Halving initially performed worse than regular Sequential Halving, but did overtake it after 25,000 rounds. Successive Rejects appeared to be doing well until the 40,000 round mark, after which it began to plateau. The results for Successive Rejects and Sequential Halving are particularly odd, flattening well before the 50,000 round limit. The reason for this is largely due to their requirement of a fixed number of arm pulls per bandit, and is discussed in greater detail below.
@@ VI-C Ludii
VI-C1 Fixed Round Number Limitations
@@
One significant drawback of the Sequential Halving and Successive Rejects algorithms is that they often do not use all of the pulls available to them [undefd, undefaa]. This is because both algorithms distribute pulls across several arm selection rounds ahead of time, requiring that the number of pulls each round be divisible between the remaining arms (which often results in some leftover pulls). This problem is particularly prevalent in situations where the number of available pulls is relatively low compared to the number of bandit arms. Our Ludii dataset presents just such a scenario, with 1085 games and a budget of 50,000 trials meaning that each game is assigned only 46 trials. Given that each game can have up to 29 potential agents that need evaluating, this very small trials-to-agents ratio severely limits the effectiveness of these approaches. Sequential Halving is most impacted by this, with only 26,834 ($\sim$54%) of its allocated trials being properly used. After this point, any remaining pulls are distributed randomly across all available arms, resulting in a substantially decreased improvement rate compare to previous rounds. Successive Rejects is also affected by this same issue, using just 39,102 ($\sim$78%) of its available trials before resorting to random pulling, but the fact that it discards its arms one at a time instead of removing half appears to reduce this impact. This issue is also much less prevalent for the GVGAI dataset, likely due to the reduced number of games this dataset has, leading to an increased number of trials per game-agent pair.
VI-D Average Probability of Error
While the primary focus of our RCP algorithm is on minimising the average simple regret across all games, we can also measure the number of games where we correctly identify the best performing agent (average probability of error). The average probability of error of each best arm identification algorithm for the GVGAI (all levels), GVGAI (level 1) and Ludii datasets are shown in Figures 6, 7 and 8 respectively.
Figure 6: Average probability of error for each Best Arm Identification algorithm when applied to the GVGAI (all levels) game-agent results dataset.
Figure 7: Average probability of error for each Best Arm Identification algorithm when applied to the GVGAI (level 1) game-agent results dataset.
Figure 8: Average probability of error for each Best Arm Identification algorithm when applied to the Ludii game-agent results dataset.
While our RCP algorithm does still outperform all other alternatives, the percentage improvement is substantially lower when measuring average probability of error (APE) compared to average simple regret. Compared to the previous second-best approach (UCB-E) RCP achieved a 10.4% lower APE for the GVGAI (all levels) dataset, a 20.2% lower APE for the GVGAI (level 1) dataset, and a 10.7% lower APE for the Ludii dataset. While this performance improvement is less substantial compared to our results for average simple regret, it is still nonetheless a demonstrable reduction in APE over prior techniques.
VI-E Uniform Bandit Selection
To better understand the impact that our dynamic multi-bandit selection process has on the overall performance of our RCP algorithm, we also evaluated our previously described uniform bandit selection variant (Section IV-B). The intention for RCP is to focus pulls on bandits where there is still a high potential for regret change, typically associated with a high uncertainty in estimating the expected reward for several top-performing arm candidates. This is in contrast to the uniform bandit selection variant, which allocates an equal number of pulls to each bandit.
The regular RCP algorithm selects the arm with the highest potential of altering the regret of it’s associated bandit, often leading to a highly imbalanced number of pulls for each bandit, see Figures 9 and 10. These graphs visualise the selection distribution of bandits (i.e. games) when applying the regular RCP algorithm (Wilson score, $\alpha$=0.05, averaged across all 10 runs) for each of our domains. We can see that RCP selects fewer trials for the majority of games when compared to uniform bandit selection. However, there are also several games with a comparatively large number of trails, indicating that these games had a relatively high regret change potential.
Figure 9: Distribution of arm pulls (i.e., number of trials) using our RCP algorithm (Wilson score, $\alpha$=0.05) for each game in the GVGAI dataset. Uniform bandit selection allocates a fixed 462 trials per game.
Figure 10: Distribution of arm pulls (i.e., number of trials) using our RCP algorithm (Wilson score, $\alpha$=0.05) for each game in the Ludii dataset. Uniform bandit selection allocates a fixed 46 trials per game.
Looking first at the average simple regret results, see Figure 11, we can see that the uniform bandit selection variant appears to achieve a better regret reduction when we have a small number of rounds, but that our RCP’s dynamic approach to selecting which bandits to pull leads to a better long-term performance boost. This most prevalent for Ludii dataset, where it achieves an average regret 17.6% smaller that it’s uniform bandit selection variant (averaged across the full 50,000 rounds).
Figure 11: Average simple regret for RCP (Wilson score, $\alpha$=0.05) when applied to each game-agent results dataset, using both dynamic (i.e., regular) and uniform bandit selection. The shaded uncertainty region indicates plus/minus one standard deviation.
In contrast, the uniform bandit selection variant seemingly performs much better when comparing the APE between the two approaches, see Figure 12. This is not unexpected, as our proposed RCP approach is designed specifically for regret reduction rather than APE, and so is likely to under-sample bandits where the selected arm is close to optimal (even if it may not actually be the best arm). However, if the goal is indeed to find the true best performing agent for each game, regardless of how small the regret difference is, then the uniform variant of RCP appears to give the best results.
Figure 12: Average probability of error for RCP (Wilson score, $\alpha$=0.05) when applied to each game-agent results dataset, using both dynamic (i.e., regular) and uniform bandit selection. The shaded uncertainty region indicates plus/minus one standard deviation.
VI-F Discussion
Overall, the improvements in best agent identification offered by our proposed RCP algorithm are clear across both domains, being able to to consistently outperform all other techniques for the vast majority of rounds tested. We believe that one of the core reasons for this improvement is our decision to focus primarily on the simple regret metric, rather than the average probability of error (APE) metric. Many prior best-arm identification algorithms choose to focus on the later, resulting in approaches that would prioritise pulling arms in a bandit that were very near the top performing arm in a attempt to find the “one true” best arm. This would often result in behaviour that provides minimal benefit towards optimising simple regret. For example, in a bandit where multiple arms achieve a very high (but not perfect) success rate, it takes a lot of pulls to identify which is the best performing arm. However, from a regret metric perspective, it makes very little difference which arm we select, and our pulls would likely be more useful on other bandits.